OpenStudy (gabbyalicorn):

Janet pays $350 annually for her gym membership plus $8 each time she takes a pilates class. Last year, Janet took 20 pilates classes. What percent of her total gym cost was represented by her pilates classes? Round the answer to the nearest tenth of a percent. A. 2.3% B. 5.7% C. 31.4% D. 45.7%

OpenStudy (gabbyalicorn):

@gaos

OpenStudy (anonymous):

Ok first we have to find the amount of money that she paid overall 350 + 8(20) = 350 + 160 = 510 She spent $160 on pilates classes and $510 overall so the percentage is 160/510 * 100% = 16/51 * 100% =31.37%

OpenStudy (anonymous):

that means that answer is C

OpenStudy (gabbyalicorn):

You are right, I double checked...Thank you!

OpenStudy (anonymous):

your welcome